OpenStudy (anonymous):

simplify (a) ^3sqrta*^4sqrta/sqrta, (b) (a^3/b^2)^5

OpenStudy (lgbasallote):

can you draw those?

OpenStudy (anonymous):

they are indices thats 3square root(a) times 4square root(a) divided by square root (a)

OpenStudy (lgbasallote):

\[\huge \sqrt [3]a \times \sqrt[4] a \div \sqrt a\] ??

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

rationals i should say

OpenStudy (lgbasallote):

\[\large \implies a^{1/3} \times a^{1/4} \div a^{1/2}\] \[\large\implies a^{4/12} \times a^{3/12} \div a^{6/12}\] \[\large \implies a^{4/12 + 3/12} \div a^{6/12}\] \[\large \implies a^{7/12} \div a^{6/12}\] \[\large\implies a^{7/12 - 6/12}\] \[\large \implies a^{1/12}\] does that help?

OpenStudy (anonymous):

thank you i see where i went wrong by not raising sqrta to the 1/2, thanks again igbasallote i have the second one under control now to

OpenStudy (lgbasallote):

welcome